A particle is acted upon by three forces equal to 50 N, 100 N and 
130 N, along the three sides of an equilateral triangle, taken in 
order. Find graphically the magnitude and direction of the resultant 
force

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Horizontal force = -50 cos60 + 100 - 130Cos60 
= 100 - 180/2 = 10N 
Vertical force = 130 Sin60 - 50 Sin60 
= 80 sqrt(3)/2 = 40 sqrt(3)
Resultant force = 10i + 4sqrt(3) j 
Hope this helps 
Good Luck 
Ajay
one year ago
 A partica is acted upon by three force eqwanet to 5N .10N and 30N along three sides of an eqwale exuregal trengal teum in find